OpenStudy (mitu12):

the perimeter of a rectangle is 52. the length is 6 greater than the width . find the length and the width of this rectangle.

OpenStudy (anonymous):

First, write an equation. Let x= width and x+6= length (the length is 6 greater than the width). Hint- Perimeter=2L+2W

OpenStudy (mitu12):

ok

OpenStudy (mitu12):

im not sure

OpenStudy (anonymous):

ok so, if the width is x, the length is x+6, and the perimeter is 52, do 2(x)+2(x+16)=52 since 2L+2W=52. Now, just solve for x. that is your width. to find your length, plug in x+6 with the number that you got for the width. that is your length

OpenStudy (mitu12):

x=5

OpenStudy (anonymous):

how did you get x=5?

OpenStudy (mitu12):

i solved the equation

OpenStudy (mitu12):

is the width 5

OpenStudy (mitu12):

and the length 30

OpenStudy (anonymous):

well if you solved the equation, you had to distribute. 2(x)+2(x+6) to get 2x+2x+12=52. Then, this becomes 4x+12=52

OpenStudy (anonymous):

solve 4x+12=52 to find x. It is not 5.

OpenStudy (mitu12):

your first reply said 16

OpenStudy (anonymous):

?

OpenStudy (mitu12):

the reply under "im not sure"

OpenStudy (anonymous):

oh mb, typo -.- sorry. anyways 4x+12=52 4x=40 x=10. x+6= 16 Width= 10 Length= 16 --------------- Check 2(10)+2(16)=52 20+32=52 52=52

OpenStudy (anonymous):

Therefore, the width is 10 and the length is 16.

OpenStudy (mitu12):

i reworked it and got those answers

OpenStudy (anonymous):

:) sorry about the typo

OpenStudy (mitu12):

no problem